value: 10.00 points A 11-year US. Treasury bond with a ace value o S1 000 pays a coupon o 575% 2.875% offace value every six months. The reported ye dto matur trs AsasX-month discount rate。5AI 2.7%) (Do not round intermediate calculations. Round your answers to 2 decimal places.) a. What is the present value of the bond? Present value b. If the yield to maturity changes to 1%, what will be the present value? Present value C. If the yield to maturity changes to 8%, what will be the present value? Present value d. if the yield to maturity changes to 15%, what will be the present value? Present value References
Answer #1
a Present value $ 1,028.75 =PV(2.7%,22,28.75,1000)*-1
b Present value $ 1,493.62 =PV(0.5%,22,28.75,1000)*-1
c Present value $     837.42 =PV(4%,22,28.75,1000)*-1
d Present value $     508.95 =PV(7.5%,22,28.75,1000)*-1